# Topological Strings in SU(3) Gauge Theory at Finite Temperature

### v1

@@ IV Numerical Setup and Results
For the simulations, we have used the lattice of size $N_{x}=N_{y}=60$, $N_{z}=4$ and $N_{\tau}=2$. The string is aligned along the $z$-direction. We consider $N_{z}=4$ as it helps suppress the fluctuations in the string length. For $N_{\tau}=2$, critical coupling for the CD transition is found to be $\beta_{c}=5.099$. Since string configurations exists only in the deconfined phase, we consider $\beta>\beta_{c}$ in our simulations.
@@
For the free energy calculations, we carry out simulations in the range $\beta_{c}-15.0$. In these simulations, thermal configurations of link variables are generated using the Cabibbo-Marinari algorithm Cabibbo:1982zn . Each link variable is updated using this heat-bath algorithm, which constitute a single sweep. Since a new configuration is generated from a previous one, there is always a non-zero autocorrelation between them. To reduce this correlation, we perform five heat-bath sweeps between successive measurements. For each $\beta$, we sample $5000$ configurations to calculate physical observables.
Figure 4: A typical string attached to $Z_{3}$ domain walls for $\beta=7.0$. The absolute value of the Polyakov loop as a function of (x,y)(top), A vector plot with the real and imaginary part of the Polyakov loop(bottom).
@@
As previously discussed, in effective models, in the presence of domain walls, the manifold $M$ effectively takes the form of a deformed triangle, as shown in Fig.3. To establish that the string configuration remains topological even in the exact theory, it is important to show that $M$ is a deformed triangle of non-zero size in the complex Polyakov loop plane. The size of $M$ is determined by the bulk equilibrium value of the Polyakov loop(largest) and the value at the core of the domain walls(smallest). As long as the smallest value remains non-zero, the size of $M$ will be non-zero, and the corresponding string configurations will be topological. We therefore calculate the magnitude of the Polyakov loop at the core of the domain walls to confirm that the size of $M$ is non-zero. The results for the Polyakov loop value at the center of the domain wall for different $\beta$ are shown in Fig.5. These results show that the Polyakov loop at the core of domain walls is indeed non-zero. Thus, the $Z_{3}$ strings resulting from the junctions of these walls are topological, as they correspond to non-trivial mappings from loops in physical space to $M$. Consequently, we observe that the Polyakov loop vanishes at the core of the $Z_{3}$ strings, see Fig.4.
Figure 5: The interface wall height vsvs $\beta$
@@
The calculation of the free energy of the string configuration requires simulations both with and without the string. This calculation involves determining the difference in action between the two cases, $\Delta S$. In the simulations with the string, the temporal links are set according to $n_{4}$. For $n_{4}=1$, all the temporal links are set to $U_{4}(n)=\mathds{1}$. For $n_{4}=2$, they are set according to the azimuthal angle coordinate($\theta_{a}$) corresponding to the position vector, $\vec{n}$. $U_{4}(n)=\mathds{1}$ for $\pi/3\leq\theta_{a}\leq\pi/3$ and $U_{4}(n)=z^{-1}(z)$ for $\pi(\pi/3)<\theta_{a}<-\pi/3(\pi)$, where $z=\exp(2\pi i/3)$. Keeping the temporal links at the boundary fixed ensures that the initial configuration thermalizes into a string with domain walls attached. A typical profile of the string at $\beta=7.0$ is shown in Fig.4, where the Polyakov loop has been averaged over along the $z$-direction. The effect of this boundary condition on the Polyakov loop profile decays within few lattice points from the boundary wall, as can be seen in Fig.4. To eliminate boundary effects, we exclude the region near the boundary where physical observables deviate from their bulk equilibrium values.
Figure 6: $\Delta S/A\beta$ vs. $\beta$ for the $Z_{3}$ interface with fitted curve
Figure 7: $\beta$ dependence of $\alpha/T^{3}$ plot
@@
In $SU(N)$ gauge theories for $N>2$, when the $Z_{N}$ symmetry is spontaneously broken, the string configuration is always attached to $N$ domain walls. The interface tension of these walls contributes significantly to the total string tension. Therefore, before discussing the string configuration, we first present our results for the interface tension of the domain walls connected to the string configurations.
@@
We calculate the interface tension of the walls away from the string core. This is because the magnitude of the Polyakov loop vanishes at the core (Fig.4), causing the domain walls to deform in that region. To avoid the effects of the core, we consider the action difference $\Delta S$, within radial annular regions around the string core, defined by $r_{a}\leq r/a\leq r_{b}$, with $r_{a}=10$ and $r_{b}=15,20$. We then compute the action difference, divide it by three (corresponding to the three interfaces), and normalize by the area of the annular patches as well as the area of the domain walls. The plot for the interface action difference with $\beta$ is shown in Fig.6. In the plot, $\Delta S_{1}$ and $\Delta S_{2}$ correspond to $r_{b}=20$ and $15$ respectively. The figure shows that within errors $\Delta S_{1}$ and $\Delta S_{2}$ agree with each other, which indicates that the choice of $r_{a}=10$ eliminates the effects of the string core. We fit the data for $\Delta S_{1,2}$ using the function $f(\beta)=e1+b1\exp(-c1(\beta-d1))+b2\exp(-c2(\beta-d2))$, which we then integrate to calculate the free energy of the interface configuration.
@@
For $\beta$ close to $\beta_{c}$, there are a lot of fluctuations in the location of the string core, consequently the action difference has larger errors. Nevertheless, we obtain a reasonable fit to the data very close to $\beta_{c}$. Fig. 7 shows our results for the interface tension, $\alpha/T^{3}$ as a function of $\beta$. These results are in good agreement with previous work Kajantie:1990bu . The consistency between the interface tension extracted from the string configuration and earlier findings confirms the reliability of our method to calculate the string tension.
@@
It is well known that the pure $SU(3)$ CD transition is weakly first-order. At the critical coupling $\beta=\beta_{c}$, the confined state ($L=0$) state and the three ($L\neq 0$) $Z_{3}$ states are degenerate. As a result, in addition to $Z_{3}$ domain walls, CD interfaces also emerge. Effective model calculations indicate that near the transition point, these domain walls decay into pairs of CD interfaces. This is also observed in the lattice simulations, where thermal fluctuations lead to the decay of domain walls into CD interfaces. Thus we expect that at the transition point, $\beta=\beta_{c}$, the interface tension to be about twice that of the CD interface.
@@
Figure 8: $\Delta S/\beta$ vs. $\beta$ for different radial patches $r/a=5,10,15,20$ fitted curve with function $f(x)$.
We now discuss the results for the free energy of the string configurations. We computed the action difference, $\Delta S$, corresponding to the various radial patches($r/a<r_{a}$), for different $\beta$ values presented in Fig. 8. These results were also fitted using the function $f(\beta)$, which was previously defined. Note that for smaller patches, as $\beta$ approaches from above, the rise in $\Delta S$ stops, and it decreases to a smaller non-zero value at $\beta_{c}$. However, this change in $\Delta S$ vs $\beta$ is in a very small range close to $\beta_{c}$, thus does not affect our results qualitatively. Following this, we perform the integration of $f(\beta)$, Eq.11, and obtain the tension or the free energy of the string. The results for $\sigma/T^{2}$ vs $\beta$, for different radial patches are shown in Fig. 9.
@@
Figure 9: $\beta$ dependence of $\sigma/T^{2}$ plot for different radial patches $r/a=5,10,15,20$.

### v2

@@ IV Numerical Setup and Results
For the simulations, we use lattices of size $N_{x}=N_{y}=60$, $N_{z}=4$ with $N_{\tau}=2,4$. The string is aligned along the $z$-direction. We consider $N_{z}=4$ as it suppresses fluctuations in the string length. For $N_{\tau}=2$ and $4$, the critical couplings for the CD transition are $\beta_{c}=5.099$ and $5.6925$, respectively. Since string configurations exist only in the deconfined phase, we consider $\beta>\beta_{c}$ in our simulations.
Table 1: Interface action difference for $N_{\tau}=2$.
$\beta$ patch 1 patch 2
$\Delta S_{1}$ error $\Delta S_{2}$ error
5.20 2.60849 0.03857 2.55569 0.02348
5.30 2.05266 0.03285 2.01550 0.02023
5.40 1.78793 0.03127 1.75478 0.01932
5.50 1.52388 0.02907 1.51606 0.01782
5.60 1.32463 0.02663 1.32239 0.01620
5.70 1.11286 0.02590 1.12360 0.01620
5.80 0.95910 0.02502 0.96622 0.01540
5.90 0.86711 0.02341 0.85801 0.01430
6.00 0.76401 0.02248 0.77086 0.01371
6.10 0.68937 0.02193 0.69040 0.01315
6.20 0.64199 0.02100 0.62903 0.01289
6.30 0.55285 0.02026 0.56255 0.01237
6.40 0.53666 0.01950 0.54253 0.01197
7.00 0.43201 0.01709 0.43442 0.01009
8.00 0.30046 0.01422 0.29513 0.00844
9.00 0.30913 0.01227 0.31334 0.00728
10.00 0.28124 0.01085 0.27210 0.00643
11.00 0.25258 0.00971 0.25486 0.00575
12.00 0.23825 0.00889 0.24045 0.00521
13.00 0.23209 0.00808 0.22244 0.00477
14.00 0.21455 0.00745 0.21222 0.00445
15.00 0.20374 0.00696 0.19966 0.00403
Table 2: Interface action difference for $N_{\tau}=4$.
$\beta$ patch 1 patch 2
$\Delta S_{1}$ error $\Delta S_{2}$ error
5.80 0.22362 0.02721 0.26723 0.01704
5.90 0.25909 0.02521 0.28556 0.01566
6.00 0.22189 0.02370 0.23457 0.01473
6.10 0.22350 0.02229 0.21348 0.01389
6.20 0.17070 0.02189 0.16103 0.01359
6.30 0.16071 0.02124 0.15917 0.01315
6.40 0.14759 0.02051 0.15209 0.01284
7.00 0.10863 0.01755 0.11986 0.01088
8.00 0.08822 0.01492 0.08180 0.00925
9.00 0.08280 0.01282 0.09236 0.00801
10.00 0.07491 0.01134 0.07590 0.00701
11.00 0.06360 0.01006 0.07088 0.00622
12.00 0.05636 0.00936 0.06180 0.00576
@@
For the free energy calculations, we carry out simulations with $\beta$ range $[5.2,15.0]$ for $N_{\tau}=2$ and $[5.8,12.0]$ for $N_{\tau}=4$. Thermal configurations of link variables are generated using the Cabibbo-Marinari algorithm [64]. Each link variable is updated using this heat-bath algorithm, which constitutes a single sweep. Since a new configuration is generated from a previous one, there is always a non-zero autocorrelation between them. To reduce this correlation, we perform five heat-bath sweeps between successive measurements. For each $\beta$, we sample 5000 configurations for $N_{\tau}=2$ and 10000 for $N_{\tau}=4$ to compute physical observables; the larger sample size for $N_{\tau}=4$ is required due to stronger fluctuations.
Figure 4: The interface wall height vsvs $\beta$ for $N_{\tau}=2$
@@
As discussed earlier, in effective models, in the presence of domain walls the manifold $M$ takes the form of a deformed triangle (Fig.3). To establish that the string configuration remains topological in the exact theory, it is necessary to show that $M$ is a deformed triangle of non-zero size in the complex Polyakov loop plane. The size of $M$ is determined by the bulk equilibrium value of the Polyakov loop (largest) and the value at the core of the domain walls (smallest). As long as the smallest value remains non-zero, the size of $M$ is non-zero and the corresponding string configurations are topological.
Figure 5: A typical string attached to $Z_{3}$ domain walls for $\beta=7.0$ for $N_{\tau}=2$. The absolute value of the Polyakov loop as a function of (x,y)(top), A vector plot with the real and imaginary part of the Polyakov loop(bottom).
Figure 6: A typical string attached to $Z_{3}$ domain walls for $\beta=10.0$ for $N_{\tau}=4$. The absolute value of the Polyakov loop as a function of (x,y)(top), A vector plot with the real and imaginary part of the Polyakov loop(bottom).
We therefore calculate the magnitude of the Polyakov loop at the core of the domain walls. The results for the Polyakov loop value at the center of the domain wall for different $\beta$ are shown for $N_{\tau}=2$ in Fig.4. We mention here that the Polyakov loop magnitude, at the core of the domain walls, decreases with decrease in temperature faster than the thermal average of the Polyakov loop. This faster decrease especially near $T_{c}$ could be due to perfect wetting, i.e nucleation of confinement-deconfinement interfaces. The results of Fig.4 show that the Polyakov loop at the core of domain walls is indeed non-zero. Thus the $Z_{3}$ strings resulting from the junctions of these walls are topological, as they correspond to non-trivial mappings from loops in physical space to $M$. Consequently, we observe that the Polyakov loop vanishes at the core of the $Z_{3}$ strings, see Fig.5 and Fig.6.
@@
The calculation of the free energy of the string configuration requires simulations both with and without the string. This involves determining the difference in action between the two cases, $\Delta S$. In the simulations with the string, the temporal links are set according to $n_{4}$. For $n_{4}<N_{\tau}$, all the temporal links are set to $U_{4}(n)=\mathds{1}$. For $n_{4}=N_{\tau}$, they are set according to the azimuthal angle coordinate ($\theta_{a}$) corresponding to the position vector $\vec{n}$. $U_{4}(n)=\mathds{1}$ for $\pi/3\leq\theta_{a}\leq\pi/3$ and $U_{4}(n)=z^{-1}(z)$ for $\pi(\pi/3)<\theta_{a}<-\pi/3(\pi)$, where $z=\exp(2\pi i/3)$.
Figure 7: $\Delta S/A\beta$ vs. $\beta$ for the $Z_{3}$ interface with fitted curve, $N_{\tau}=2$
Figure 8: $\beta$ dependence of $\alpha/T^{3}$ plot for $N_{\tau}=2$
Figure 9: $\Delta S/A\beta$ vs. $\beta$ for the $Z_{3}$ interface with fitted curve, $N_{\tau}=4$
Figure 10: $\beta$ dependence of $\alpha/T^{3}$ plot for $N_{\tau}=4$
Keeping the temporal links at the boundary fixed ensures that the initial configuration thermalizes into a string with domain walls attached. Further, the boundary conditions ensures that the average angular separation between the domain walls is $2\pi/3$, corresponding to the lowest free energy configurations. A typical profile of the string at $\beta=7.0$ for $N_{\tau}=2$ and at $\beta=10.0$ for $N_{\tau}=4$ is shown in Fig.5 and Fig.6 respectively, where the Polyakov loop is averaged over the $z$-direction. The string configuration shows that the domain walls broaden near the string junction. Our observations indicate that this is due to presence of other domain walls at the junction. This broadening resembles the broadening observed close to the confinement-deconfinement transition due to perfect wetting. The effect of this boundary condition on the magnitude profile of the Polyakov loop, decays within a few lattice points from the boundary wall. To eliminate boundary effects, we exclude the region near the boundary where physical observables deviate from their bulk equilibrium values.
@@
In $SU(N)$ gauge theories with $N>2$, when the $Z_{N}$ symmetry is spontaneously broken, the string configuration is always attached to $N$ domain walls. The interface tension of these walls contributes significantly to the total string tension. Therefore before discussing the string configuration, we first present our results for the interface tension of the domain walls connected to the string configurations.
@@
We calculate the interface tension away from the string core. This is because the magnitude of the Polyakov loop vanishes at the core (Fig.5 and Fig.6), causing the domain walls to deform in that region. To avoid these effects we consider the action difference $\Delta S$ within radial annular regions around the string core, defined by $r_{a}\leq r/a\leq r_{b}$, with $r_{a}=10$ and $r_{b}=15,20$. We then compute the action difference, divide it by three (corresponding to the three interfaces), and normalize by the area of the annular patches as well as the area of the domain walls. The data for the interface action difference are presented in Tables 1 and 2. The corresponding plots as a function of $\beta$ are shown in Figs. 7 and 9 for $N_{\tau}=2$ and $4$, respectively. In the tables, the results for the patch with $r_{b}=15$($r_{b}=20$) are referred to as patch 1(patch 2). In the plot, $\Delta S_{1}$ and $\Delta S_{2}$ correspond to $r_{b}=20$ and $15$ respectively. The weak coupling analysis suggests that for large $\beta$ the action difference is expected to behave as $\sim 1/\sqrt{\beta}$, [45]. We find that our measured values of action differences, are compatible with the weak coupling results. However, to fit the data in the whole range of $\beta$, we consider the following function $f(\beta)=b_{1}\exp[-c_{1}(\beta-d_{1})]+b_{2}\exp[-c_{2}(\beta-d_{2})]+e_{1}$. We then integrate this fitted function to calculate the free energy of the interface configuration.
@@
For $\beta$ close to $\beta_{c}$ there are large fluctuations in the location of the string core, leading to larger errors in the action difference. Nevertheless, we obtain a reasonable fit to the data very close to $\beta_{c}$. The results for the interface tension $\alpha/T^{3}$ as a function of $\beta$ are shown in Fig.8 and Fig.10 corresponding to $N_{\tau}=2$ and $4$. The results for $N_{\tau}=2$ qualitatively agree with previous work [32], confirming the reliability of our method.
@@
It is well known that the pure $SU(3)$ CD transition is weakly first-order. At $\beta=\beta_{c}$ the confined state ($L=0$) and the three ($L\neq 0$) $Z_{3}$ states are degenerate. As a result, in addition to $Z_{3}$ domain walls, CD interfaces also emerge. Effective model calculations indicate that near the transition point these domain walls decay into pairs of CD interfaces. This behaviour is also observed in the lattice simulations, where due to perfect wetting locally CD interfaces can nucleate near the domain wall core, effectively broadening the domain walls and eventually, for temperatures near $T_{c}$, lead to the decay of domain walls into CD interfaces, [38, 39]. Thus at $\beta=\beta_{c}$ the interface tension is expected to be approximately twice that of the CD interface. In the following we present our results for the string tension for $N_{\tau}=2$ and $N_{\tau}=4$.
IV.1 $N_{\tau}=2$ results
For $N_{\tau}=2$, we compute the action difference $\Delta S$ corresponding to different radial patches around the string core. The results for $\Delta S/\beta$ as a function of $\beta$ are shown in Fig.11 and the corresponding data is given in Table 3. We mention here that the action difference will constitute the contribution from the domain walls as well as the string core. Unlike the contribution from the domain walls, there are no known results for the string core in the weak coupling limit. However, reasonable fit to the data for the whole range of $\beta$ considered for simulations was achieved with the function $f(\beta)$ defined earlier. For smaller patches, as $\beta$ approaches $\beta_{c}$ from above, the rise in $\Delta S$ stops and decreases to a smaller non-zero value at $\beta_{c}$. However, this change occurs in a very narrow region close to $\beta_{c}$ and does not affect the results qualitatively.
@@ IV.1 $N_{\tau}=2$ results
Figure 11: For $N_{\tau}=2$, $\Delta S/\beta$ vs. $\beta$ for different radial patches $r/a=5,10,15,20$ fitted curve with function $f(x)$.
@@
Figure 12: For $N_{\tau}=2$, $\beta$ dependence of $\sigma/T^{2}$ plot for different radial patches $r/a=5,10,15,20$.
Following this, we integrate $f(\beta)$ using Eq.7 to obtain the free energy of the string configuration. The resulting string tension $\sigma/T^{2}$ as a function of $\beta$ for different radial patches is shown in Fig.12. A steep rise in $\sigma/T^{2}$ is observed close to $\beta_{c}$, followed by an approximately linear increase for larger $\beta$.
Table 3: Action difference $\Delta S$ and jackknife error for $N_{\tau}=2$
$\beta$ $r/a=5$ $r/a=10$ $r/a=15$ $r/a=20$
$\Delta S$ errorerror $\Delta S$ errorerror $\Delta S$ errorerror $\Delta S$ errorerror
5.2 110.36484 0.60116 257.87405 1.25330 402.49983 1.86822 539.88657 0.41658
5.3 103.59388 0.54066 224.67773 1.07956 341.85150 1.61047 453.54355 0.38919
5.4 92.62027 0.50698 197.82285 1.00640 300.20597 1.50576 401.59460 0.32945
5.5 82.39651 0.47625 174.47203 0.95316 263.79369 1.39979 353.10258 0.30470
5.6 71.70216 0.45873 149.49289 0.90595 226.07948 1.33539 303.65829 0.28723
5.7 61.30492 0.43776 127.88278 0.86314 194.87043 1.26551 262.55875 0.27555
5.8 53.40427 0.40917 111.89299 0.81756 169.56406 1.20557 226.53799 0.26649
5.9 47.13730 0.40463 99.47629 0.79970 150.55162 1.15040 201.96085 0.25441
6.0 41.00473 0.37837 88.08466 0.76249 132.83176 1.12734 179.68027 0.24403
6.1 36.47997 0.37304 78.15178 0.72854 119.75076 1.10030 161.14301 0.23580
6.2 33.01000 0.36098 72.43353 0.70372 109.79351 1.05420 147.18480 0.22791
6.3 31.43271 0.34900 66.88779 0.69024 100.48857 1.02028 134.15992 0.22342
6.4 29.26770 0.34022 62.58853 0.66935 94.97354 0.98328 128.83123 0.21129
7.0 20.97773 0.28890 47.49854 0.58267 72.86676 0.84560 98.35134 0.18141
8.0 16.62055 0.24176 35.60965 0.48075 53.88290 0.70482 71.32716 0.15119
9.0 14.75837 0.21165 33.48775 0.41564 51.24049 0.60717 68.67947 0.12853
10.0 13.04334 0.18806 29.76977 0.37084 46.83962 0.53784 62.64910 0.11453
11.0 11.74472 0.17133 27.68351 0.33805 42.57028 0.49007 57.46369 0.10168
12.0 10.89854 0.15542 25.42321 0.29672 39.76455 0.43152 54.15185 0.09198
13.0 9.85393 0.14322 23.71855 0.27401 37.22653 0.39858 50.06809 0.08413
14.0 9.29583 0.13142 22.70685 0.25704 35.26673 0.37050 48.20444 0.07886
15.0 8.54539 0.11975 21.33430 0.23581 33.56463 0.34129 45.18588 0.07070
Table 4: Action difference $\Delta S$ and jackknife error for $N_{\tau}=4$
$\beta$ $r/a=5$ $r/a=10$ $r/a=15$ $r/a=20$
$\Delta S$ errorerror $\Delta S$ errorerror $\Delta S$ errorerror $\Delta S$ errorerror
5.8 5.68990 1.19789 15.59788 2.50520 29.01490 4.01806 47.66482 5.53609
5.9 7.31175 0.89521 19.02333 1.65565 34.56860 2.45648 53.29115 3.51277
6.0 5.91528 0.76407 16.92231 1.54608 30.23584 2.45368 45.07127 3.11033
6.1 5.58601 0.72473 14.89238 1.40068 28.30254 2.10851 40.51041 2.61986
6.2 6.07489 0.64556 15.42404 1.29691 25.66628 1.98814 34.74756 2.53741
6.3 5.56974 0.57049 15.55875 1.17513 25.20144 1.86186 34.65953 2.21769
6.4 4.78848 0.52888 13.67205 1.16905 22.52737 1.81993 31.92254 2.50208
7.0 4.38683 0.55541 10.79059 1.03709 17.30854 1.51373 25.17335 1.92691
8.0 2.91041 0.43509 9.39564 0.83151 14.68870 1.19849 19.21144 1.65528
9.0 2.28381 0.36117 7.60547 0.72997 12.57344 1.13864 18.68845 1.59952
10.0 2.39689 0.33818 6.36685 0.64099 10.86126 1.05836 15.47505 1.37010
11.0 2.02822 0.28034 6.00958 0.54015 9.82546 0.84510 14.51481 1.09442
12.0 1.59492 0.24851 4.90090 0.48042 8.28244 0.75817 12.31742 1.04125
IV.2 $N_{\tau}=4$ results
Figure 13: For $N_{\tau}=4$, $\Delta S/\beta$ vs. $\beta$ for different radial patches $r/a=5,10,15,20$ fitted curve with function $f(x)$.
We repeat the same analysis for $N_{\tau}=4$ lattices. The action difference $\Delta S$ is calculated for the radial patches around the string core Fig.13 and fitted using the same function $f(\beta)$ and the corresponding data is given in Table 4. The qualitative behaviour of $\Delta S$ as a function of $\beta$ remains similar to that observed for $N_{\tau}=2$. Near $\beta_{c}$ the fluctuations in the string core position again lead to larger statistical errors, while for larger $\beta$ the behaviour becomes smoother.
Integrating the fitted function $f(\beta)$ yields the free energy of the string configuration and the corresponding string tension. The dependence of $\sigma/T^{2}$ on $\beta$ shown in Fig.14. We see that the free energy for the string as well as the domain walls rises slower near $\beta_{c}$ as compared to $N_{\tau}=2$, with a rapid rise near $\beta_{c}$ and an approximately linear increase at larger $\beta$. These results indicate that the qualitative features of the string configurations remain unchanged with the lattice cut-off.
Figure 14: For $N_{\tau}=4$, $\beta$ dependence of $\sigma/T^{2}$ plot for different radial patches $r/a=5,10,15,20$.